OpenStudy (anonymous):

the rectangular pen for a pet is 6 feet longer athan it is wide. Give possible values for the width W of the pen if its area must be between 187 and 520 square feet, inclusively.

OpenStudy (anonymous):

is a start \[187\leq W(W+6)\leq520\]

OpenStudy (anonymous):

do you know how to go from here? you are going to have to solve two inequalities \[w(w+6)\leq 520\] \[w^2+6w\leq 520\] \[w^2+6w-520\leq 0\] \[(w-20)(w+26)\leq 0\] so \[-26\leq w \leq 20\]

OpenStudy (anonymous):

then you have to solve \[187\leq w^2+6w\] \[0\leq w^2+6w-187\] \[0\leq(w-11)(w+17)\] so either \[w\leq -17\] or \[w\geq 11\]

OpenStudy (anonymous):

a pain but at least both of these factored. now put the inequalities together to get \[11\leq w\leq 20\]

OpenStudy (anonymous):

thank u.......